OpenStudy (anonymous):

what is 8 plus four minus 6 one

OpenStudy (anonymous):

8+4-6=6

OpenStudy (anonymous):

8+4 + (-6)+ (-1)= X 8 + 4= 12 right? -6+-1= -7 What is 12+ (-7)? Hint:In this case, adding a negative number is the same as subtracting it from a positive number.

OpenStudy (anonymous):

The answer would be (-5) and.....